# Effective synchronization amid noise-induced chaos

### v1

@@ A.2 Identifying a fiducial phase
If one kicks an oscillator that was at $\varphi=0$ (empty squares), it will then undergo the trajectories shown with the dashed red curves in Fig. S2. After $9T,10T,11T,\ldots$, we find, to good accuracy, that it reaches the points marked by solid squares, which is found to have a phase $\psi\simeq-0.13$ for the weak forcing (a) and $\psi\simeq-0.44$ for the strong forcing (b). (As before, $\psi$ is still defined mod $\mathrm{mod}\,1$. However, so to obtain the smooth collection of blue empty points in Fig. S3, we determined integer offsets by counting the number of cycles completed during $9T$ relative to the $9$ cycles that an unperturbed oscillator would have completed during $9T$.) Likewise, kicking oscillators that are at $\varphi\simeq 0.35$ (empty triangles) and $\varphi\simeq 0.6$ (empty diamonds), through the dotted-orange and full-blue trajectories, they will, respectively, reach $\psi\simeq 0.73$ and $\psi\simeq 0.36$ for the weak forcing (a) and $\psi\simeq 1.37$ and $\psi\simeq-0.66$ for the strong forcing (b), which are depicted with solid triangles and diamonds. We remind that were the three not kicked, after $9T$ they would have returned back to their original positions (00, $0.35$, and $0.6$), and that any phases that differ by an integer necessarily correspond to the same point on the orbit.
To determine the final phase $\psi$ for arbitrary initial phases $\varphi$, we repeat this process for many closely spaced initial $\varphi$’s. The result is the phase maps shown in Fig. S3, which depend on the displacement (forcing magnitude), shown in Fig. S3. The lags between the initial and final phases are shown with the empty blue symbols in Fig. S2. The squares, triangles, and diamonds correspond to the final phase $\psi$ of $\varphi=0$, $\varphi\simeq 0.35$, and $\varphi\simeq 0.6$, respectively, under each forcing. Constructing the phase map $\psi(\varphi)$ amounts to taking mod $\mathrm{mod}\,1$ of the resulting lags. Since the system is rotation invariant (as the next kick is also randomly timed according to a uniform distribution $[0,1)$), we shift all phases such that, arbitrarily, $\psi(\varphi=0)=0$. This allows the comparison with the cubic phase map considered in the main text. Upon this shift and taking the moduli, we find the phase maps depicted with connected red points. Indeed the very strong forcing $z\to z+0.45$ leads to a very violent phase map (even having an apparent discontinuity in the spreading factor $d\psi/d\varphi$ at $\varphi\simeq 0.32$). The mild forcing $z\to z+0.15$, which is the limit with which we are concerned in this work, has given rise to a well-behaved and smooth phase map with a single minimum and maximum. As we explained in the main text, broad minima and maxima are the ones facilitating the formation of sharply-peaked and thence low-entropy distributions. This suggests that the findings of the main text for the simplistic cubic phase map we chose should apply in ‘real-world’ forcings as well.
@@
Here we revisit the discussion of Sec. III of the main text for a synchronizing phase map with $\Lambda<0$. That is, the agents’ phases agree up to a gradually-diminishing uncertainty. For $\Lambda<0$, both distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ and $q^{(k)}_{\mathrm{B}}(\varphi)$ are expected to become a sharp peak centered around the same phase value. To test this, we initiate both distributions to be concentrated uniformly into eccentric narrow sectors of angular width $u=0.05$, one centered at $\varphi=0.025$ and the other at $\varphi=0.525$. While indeed both entropies decay to $\infty$- ∞ (see Fig. S6(b)), both KLDs have reached a different nonzero plateau (see Fig. S6(a)).
We can analyze this observation quantitatively. Since each $q_{\mathrm{A}}^{(k)}(\varphi)$ and $q_{\mathrm{B}}^{(k)}(\varphi)$ is very narrow, for a given $\beta^{(k)}$, at most a single $\hat{\varphi}$ contributes appreciably to a given $\varphi$ during a kick event. Namely, if there is a $\hat{\varphi}$ such that $\psi(\hat{\varphi}+\beta^{(k)})=\varphi$ and $q_{\mathrm{A}}^{(k)}(\hat{\varphi})$ is within its peak region, then it is given by
@@
[…]
and similarly $D^{(k)}(\mathrm{B}\|\mathrm{A})=D^{(k-1)}(\mathrm{B}\|\mathrm{A})$. Thus, indeed the KLD is capable of remaining constant for two infinitely narrow distributions, as they either “stretch” or “contract” together to the same extent. Similar reasoning explains why, while the entropies decay to $\infty$- ∞, the difference among them for these two narrow distributions remains constant for low entropy values in Fig. S6(b).
@@
Based on this finding, we suggest that the effect of these large $\Delta\varphi_{\mathrm{f}}$’s can be much reduced. Each agent is able to determine the uncertainty of its highest and second highest peak. If the two heights are statistically indistinguishable, the agent cannot obtain an unambiguous $\varphi_{\mathrm{f}}$ and would not be able infer the other agent’s effective phase at that kick. Had this been done in our analysis, the large $|\Delta\varphi_{\mathrm{f}}|$’s could have been reduced for distributions whose entropies are not approaching 00. We anticipate that any remaining large $\varphi_{\mathrm{f}}$’s could be made insignificant.
These misidentifications were an important factor in our choice of smoothing width $\sigma=e^{S}$. Indeed, when we used $\sigma=e^{S}/N$ (rather than $\sigma=e^{S}$ of the main text) as a bandwidth, there were $791$ misidentified points ($8\%$), which have occurred with distributions of entropies as a low as $S=-14$; see Fig. S8(b). In fact, the large concentration of points surrounding $|\Delta\varphi_{\mathrm{f}}|=e^{S}$ suggests that the bandwidth is too small, and not enough samples are involved in estimating the peak center, raising the error from $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}/\sqrt{N}$ to just the overall “width” of a given distribution $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}$. At the same time, in the cases where the identification of the peak was precise (that is, the highest peak clearly dominates over the rest and encapsulated $\sim N$ samples), this smaller bandwidth allowed for an appropriately smaller discrepancy, $|\Delta\varphi_{\mathrm{f}}|\sim(e^{S}/N)/\sqrt{N}$. Thus, as we expected, smaller bandwidths facilitate smaller deviations, which comes at a cost of more frequent, finite-sample-size-related misidentifications of the highest peak.
@@
Figure S5: A sequence of two, initially-identical phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to different noises in a typical realization of Eq. (1) and Eq. (2) of the main text with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\sim\mathrm{U}[0.5,0.55)$. The distributions are drawn on a circle so the periodicity mod $\,\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. Since the noise is different for each ensemble, $\Lambda<0$ produces different synchronized states, whereas $\Lambda>0$ produces erratic distributions that do not converge unto each other. The distributions are sampled using $N=500$ oscillators via the nearest-neighbor distances; see Eq. (A1) of the main text. The initial samples drawn for ensemble ${\mathrm{A}}$ are copied to ensemble ${\mathrm{B}}$, so the starting point is truly identical.
@@
Figure S7: Illustration of the kernel-density estimation method. Here we estimate the distribution from $N=500$ phase samples obtained at $k=100$ by agent $\mathrm{B}$ during the realization of Fig. 2 of the main text for $\Lambda=0.141$. The density estimate of Eq. (A1) of the main text is shown in points. The kernel-density estimate, $Q(\varphi;\sigma)$, is carried out with either $\sigma=e^{S}$ (the bandwidth we use throughout Sec. IV of the main text; dashed red curve) or $\sigma=e^{S}/N$ (the alternative extreme; thin solid black curve). $S=-2.29$ for this distribution using Eq. (A2) of the main text. A scale bar of $e^{S}$ is shown for reference with a thick black horizontal line. Clearly, the result of the latter bandwidth is not statistically significant, as it is comparable to the nearest-sample distances. The analysis in the main text is carried out with the former bandwidth.

### v2

@@ A.2 Identifying a fiducial phase
If one kicks an oscillator that was at $\varphi=0$ (empty squares), it will then undergo the trajectories shown with the dashed red curves in Fig. S2. After $9T,10T,11T,\ldots$, we find, to good accuracy, that it reaches the points marked by solid squares, which is found to have a phase $\psi\simeq-0.13$ for the weak forcing (a) and $\psi\simeq-0.44$ for the strong forcing (b). (As before, $\psi$ is still defined $\mathrm{mod}\,1$. However, so to obtain the smooth collection of blue empty points in Fig. S3, we determined integer offsets by counting the number of cycles completed during $9T$ relative to the $9$ cycles that an unperturbed oscillator would have completed during $9T$.) Likewise, kicking oscillators that are at $\varphi\simeq 0.35$ (empty triangles) and $\varphi\simeq 0.6$ (empty diamonds), through the dotted-orange and full-blue trajectories, they will, respectively, reach $\psi\simeq 0.73$ and $\psi\simeq 0.36$ for the weak forcing (a) and $\psi\simeq 1.37$ and $\psi\simeq-0.66$ for the strong forcing (b), which are depicted with solid triangles and diamonds. We remind that were the three not kicked, after $9T$ they would have returned back to their original positions (0, $0.35$, and $0.6$), and that any phases that differ by an integer necessarily correspond to the same point on the orbit.
To determine the final phase $\psi$ for arbitrary initial phases $\varphi$, we repeat this process for many closely spaced initial $\varphi$’s. The result is the phase maps shown in Fig. S3, which depend on the displacement (forcing magnitude), shown in Fig. S3. The lags between the initial and final phases are shown with the empty blue symbols in Fig. S2. The squares, triangles, and diamonds correspond to the final phase $\psi$ of $\varphi=0$, $\varphi\simeq 0.35$, and $\varphi\simeq 0.6$, respectively, under each forcing. Constructing the phase map $\psi(\varphi)$ amounts to taking $\mathrm{mod}\,1$ of the resulting lags. Since the system is rotation invariant (as the next kick is also randomly timed according to a uniform distribution $[0,1)$), we shift all phases such that, arbitrarily, $\psi(\varphi=0)=0$. This allows the comparison with the cubic phase map considered in the main text. Upon this shift and taking the moduli, we find the phase maps depicted with connected red points. Indeed the very strong forcing $z\to z+0.45$ leads to a very violent phase map (even having an apparent discontinuity in the spreading factor $d\psi/d\varphi$ at $\varphi\simeq 0.32$). The mild forcing $z\to z+0.15$, which is the limit with which we are concerned in this work, has given rise to a well-behaved and smooth phase map with a single minimum and maximum. As we explained in the main text, broad minima and maxima are the ones facilitating the formation of sharply-peaked and thence low-entropy distributions. This suggests that the findings of the main text for the simplistic cubic phase map we chose should apply in ‘real-world’ forcings as well.
@@
Here we revisit the discussion of Sec. III of the main text for a synchronizing phase map with $\Lambda<0$. That is, the agents’ phases agree up to a gradually-diminishing uncertainty. For $\Lambda<0$, both distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ and $q^{(k)}_{\mathrm{B}}(\varphi)$ are expected to become a sharp peak centered around the same phase value. To test this, we initiate both distributions to be concentrated uniformly into eccentric narrow sectors of angular width $u=0.05$, one centered at $\varphi=0.025$ and the other at $\varphi=0.525$. While indeed both entropies decay to -$\infty$ (see Fig. S6(b)), both KLDs have reached a different nonzero plateau (see Fig. S6(a)).
[…]
@@
$D^{(k)}(\mathrm{A}\|\mathrm{B})=\int_{0}^{1}d\hat{\varphi}q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})\ln\frac{q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})}{q_{\mathrm{B}}^{(k-1)}(\hat{\varphi})}=D^{(k-1)}(\mathrm{A}\|\mathrm{B}),$ (S4)
and similarly $D^{(k)}(\mathrm{B}\|\mathrm{A})=D^{(k-1)}(\mathrm{B}\|\mathrm{A})$. Thus, indeed the KLD is capable of remaining constant for two infinitely narrow distributions, as they either “stretch” or “contract” together to the same extent. Similar reasoning explains why, while the entropies decay to -$\infty$, the difference among them for these two narrow distributions remains constant for low entropy values in Fig. S6(b).
@@
Based on this finding, we suggest that the effect of these large $\Delta\varphi_{\mathrm{f}}$’s can be much reduced. Each agent is able to determine the uncertainty of its highest and second highest peak. If the two heights are statistically indistinguishable, the agent cannot obtain an unambiguous $\varphi_{\mathrm{f}}$ and would not be able infer the other agent’s effective phase at that kick. Had this been done in our analysis, the large $|\Delta\varphi_{\mathrm{f}}|$’s could have been reduced for distributions whose entropies are not approaching 0. We anticipate that any remaining large $\varphi_{\mathrm{f}}$’s could be made insignificant.
These misidentifications were an important factor in our choice of smoothing width $\sigma=e^{S}$. Indeed, when we used $\sigma=e^{S}/N$ (rather than $\sigma=e^{S}$ of the main text) as a bandwidth, there were $791$ misidentified points ($8\%$), which have occurred with distributions of entropies as a low as S=14S=-14; see Fig. S8(b). In fact, the large concentration of points surrounding $|\Delta\varphi_{\mathrm{f}}|=e^{S}$ suggests that the bandwidth is too small, and not enough samples are involved in estimating the peak center, raising the error from $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}/\sqrt{N}$ to just the overall “width” of a given distribution $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}$. At the same time, in the cases where the identification of the peak was precise (that is, the highest peak clearly dominates over the rest and encapsulated $\sim N$ samples), this smaller bandwidth allowed for an appropriately smaller discrepancy, $|\Delta\varphi_{\mathrm{f}}|\sim(e^{S}/N)/\sqrt{N}$. Thus, as we expected, smaller bandwidths facilitate smaller deviations, which comes at a cost of more frequent, finite-sample-size-related misidentifications of the highest peak.
@@
Figure S5: A sequence of two, initially-identical phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to different noises in a typical realization of Eq. (1) and Eq. (2) of the main text with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\sim\mathrm{U}[0.5,0.55)$. The distributions are drawn on a circle so the periodicity $\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. Since the noise is different for each ensemble, $\Lambda<0$ produces different synchronized states, whereas $\Lambda>0$ produces erratic distributions that do not converge unto each other. The distributions are sampled using $N=500$ oscillators via the nearest-neighbor distances; see Eq. (A1) of the main text. The initial samples drawn for ensemble $\mathrm{A}$} are copied to ensemble $\mathrm{B}$}, so the starting point is truly identical.
@@
Figure S7: Illustration of the kernel-density estimation method. Here we estimate the distribution from $N=500$ phase samples obtained at $k=100$ by agent $\mathrm{B}$ during the realization of Fig. 2 of the main text for $\Lambda=0.141$. The density estimate of Eq. (A1) of the main text is shown in points. The kernel-density estimate, $Q(\varphi;\sigma)$, is carried out with either $\sigma=e^{S}$ (the bandwidth we use throughout Sec. IV of the main text; dashed red curve) or $\sigma=e^{S}/N$ (the alternative extreme; thin solid black curve). S=2.29S=-2.29 for this distribution using Eq. (A2) of the main text. A scale bar of $e^{S}$ is shown for reference with a thick black horizontal line. Clearly, the result of the latter bandwidth is not statistically significant, as it is comparable to the nearest-sample distances. The analysis in the main text is carried out with the former bandwidth.